Evaluate the exponential equation for three positive values of x, three negative values of x, and at x=0. Show your work. Use the resulting ordered pairs to plot the graph; State the equation of the line asymptotic to the graph (if any). y=5^x
Answer by agille32(1)

y=5x
y=5^-3 = .008
y=5^-2 = .04
y=5^-1 = 0.2
y=5^0 = 1
y=5^1=5 =The logarithm, base a, of a is always 1: loga a = 1.
y=5^2=5*5=25
y=5^3=5*5*5=125